# Bayesian Optimization in Language Space: An Eval-Efficient AI Self-Improvement Framework

### v2

@@ 8 Conclusion
The importance of developing evaluation-efficient self-improving AI systems is substantial for its societal applications, yet this problem has remained unidentified and unaddressed. Motivated by the effectiveness of prompt optimization, which outperforms reinforcement fine-tuning methods, we formulate this problem as an iterative prompt optimization and address it.
We propose T-BoN BO, a simple, evaluation-efficient prompt optimization framework that leverages textual edits and Best-of-N search to guide prompt improvement. We prove that T-BoN BO BO statistically emulates UCB-BO, which is a provably evaluation-efficient algorithm. Experiments simulating ad alignment across eight creative-ad tasks demonstrate that T-BoN BO outperforms strong state-of-the-art baselines, including large Best-of-N and GEPA. We validate that these improvements are not attributed to the usage of specific persona dataset or trivial improvements from using poor initial prompts.
Our paper makes three key contributions to the computer science and marketing literature. First, it identifies evaluation efficiency as an essential objective for a societal self-improving system design. Second, it presents a simple yet provably efficient AI self-improving system design framework that can be applied to real-world societal systems, e.g., ad agencies, that incur high evaluation costs. Third, we theoretically prove that the presented framework is evaluation-efficient.
In sum, this work bridges classical Bayesian optimization and prompt-based self-improving AI systems, demonstrating that evaluation-efficient optimization in the prompt space is both theoretically tractable and practically impactful. As organizations increasingly rely on AI-generated content, methods like T-BoN BO can help managers achieve higher-quality outputs without proportionally increasing evaluation burdens.
Appendix A Prompts used for digital marketing
comparison_\_prompt =
RESPOND WITH ONLY THE NUMBER 1 OR 2. NOTHING ELSE.
1 = first image is better
2 = second image is better ”
Figure 11: $M_{\text{critic}}$’s pairwise tournament prompt for digital marketing example.
VISUAL EXAMPLES – BEST VS WORST PERFORMING:
RANK 1/10 WORST PERFORMING (Score: 2.14/5.0):
Prompt excerpt: <prompt>…
[Image]
RANK 2/10 LOWER HALF (Score: 2.52/5.0):
Prompt excerpt: <prompt>…
[Image]
$\ldots$ RANK 9/10 UPPER HALF (Score: 3.81/5.0):
Prompt excerpt: <prompt>…
[Image]
RANK 10/10 BEST PERFORMING (Score: 4.15/5.0):
Prompt excerpt: <prompt>…
[Image]
Figure 12: Meta reflection prompt for digital marketing example.
@@ Appendix A Prompts used for digital marketing
Prompt used for generating prompts for ads generation Generate a creative, structured and descriptive prompt for a generative AI model (e.g., Imagen) that will produce a brand-aligned, and scroll-stopping advertisement image suitable for an Instagram feed.
@@ 5. What to include:
Figure 13: Prompt for generating the prompts for the initial 64 ads.
Prompt for generating the worst-of-64 in GreenBite (plant-based burger patty) ad scenario. **Key Message:** To visually communicate the core idea: "Experience the juicy, no-compromise satisfaction of a real burger, made better for you and the planet."
@@
Prompt for generating the best-of-64 in GreenBite (plant-based burger patty) ad scenario. **1. Key Message (The "Why"):** To visually capture the peak moment of guilt-free indulgence, proving that a plant-based burger can deliver an unbelievably juicy, satisfying, and no-compromise flavor experience identical to a classic beef burger. The text overlay will read: "The No-Compromise Burger."

### v3

@@ 10 Conclusion
Evaluation-efficient self-improving AI is increasingly important in societal and business settings where real-world feedback (e.g., human responses, field experiments) is slow and costly, even as candidate generation and analysis have become cheap. In this paper, we address this gap by proposing TextBO, a simple evaluation-efficient prompt-optimization framework that combines TextGrad-style local textual updates with Best-of-N gradient selection. Our main theoretical result shows that Best-of-N over locally sampled textual edits induces, in probability, ascent directions of a UCB-style acquisition function. This provides a principled bridge between language-space self-improvement and classical UCB Bayesian optimization: TextBO emulates parallel gradient-based UCB-BO in the implicit embedding space induced by the LLM, while operating purely in language space and without constructing an explicit surrogate or calibrated uncertainty model. Empirically, we find that TextBO improves evaluation efficiency in ad-creative prompt optimization, and our ablations indicate that the gains are not driven by a particular persona dataset or by starting from unusually weak prompts. We additionally find that the same Best‑of‑N multi‑step mechanism strengthens GEPA in agentic benchmark settings: TextBO‑GEPA exhibits larger improvements than GEPA across evaluation steps on agentic AI benchmarks.
Overall, this work makes four contributions to literature on AI, optimization, and marketing. First, we elevate evaluation efficiency as a first-order design objective for self-improving AI deployed in real-world decisions. Second, we introduce TextBO, a simple and practical language-space optimization procedure that combines textual gradients with Best-of-N selection. Third, we provide theory showing that Best-of-N textual-gradient steps implement optimism-driven UCB-style exploration in an implicit embedding space, yielding evaluation-efficiency guarantees. Fourth, we demonstrate empirical gains relative to strong baselines in both the digital advertising case study and the agentic AI benchmarks.
From a managerial perspective, evaluation-efficient self-improving AI can sharply reduce the marginal cost of iteration in data-driven decision making and enable even small businesses and resource constrained organizations to adopt AI tools (in ad optimization, promotion design, and other marketing activities). As generation and analysis become automated, our framework offers firms the tools and ability to sustain continuous optimization loops that adapt campaigns to evolving market conditions and consumer responses.
Appendix A GP-UCB and Its Evaluation Efficiency
Setup.
[…]
The objectives.
$\mathcal{R}_{C}(T;f,\mathcal{A}):=\sum_{t=1}^{T}\left[\max_{\mathbf{x}\in\mathcal{D}}f(\mathbf{x})-f\left(\mathbf{x}_{t}\right)\right]$ (A1)
The other is simple regret, defined as
$\mathcal{R}_{S}(T;f,\mathcal{A}):=\max_{\mathbf{x}\in\mathcal{D}}f(\mathbf{x})-f\left(\mathbf{x}^{(T)}\right)$ (A2)
The lower bound.
$\inf_{\mathcal{A}}\sup_{\|f\|_{\mathcal{N}_{\Psi}(\mathcal{D})}\leq B}\mathbb{E}\left[\mathcal{R}_{C}(T;f,\mathcal{A})\right]=\begin{cases}\Omega(T^{\frac{\nu+d}{2\nu+d}}),&\text{ for Matérn kernels, }\\ \Omega(T^{\frac{1}{2}}\ln^{\frac{d}{2}}(T)),&\text{ for SE kernels. }\end{cases}$ (A3)
$\inf_{\mathcal{A}}\sup_{\|f\|_{\mathcal{N}_{\Psi}(\mathcal{D})}\leq B}T\left(\epsilon,f,\mathcal{A}\right)=\begin{cases}\Omega\left(\left(\frac{1}{\epsilon}\right)^{2+d/\nu}\right),&\text{ for Matérn kernels, }\\ \Omega\left(\frac{1}{\epsilon^{2}}\left(\log\frac{1}{\epsilon}\right)^{d/2}\right),&\text{ for SE kernels. }\end{cases}$ (A4)
where Matérn kernels are defined as
$\Psi_{\mathrm{M}}\left(\mathbf{x}-\mathbf{x}^{\prime}\right):=\frac{1}{\Gamma(\nu)2^{\nu-1}}\left(\frac{2\sqrt{\nu}\left\|\mathbf{x}-\mathbf{x}^{\prime}\right\|_{2}}{\ell}\right)^{\nu}K_{\nu}\left(\frac{2\sqrt{\nu}\left\|\mathbf{x}-\mathbf{x}^{\prime}\right\|_{2}}{\ell}\right),\quad\mathbf{x},\mathbf{x}^{\prime}\in\mathcal{D}$ (A5)
and SE kernels are defined as
$\Psi_{\mathrm{SE}}\left(\mathbf{x}-\mathbf{x}^{\prime}\right)=\exp\left(-\frac{\left\|\mathbf{x}-\mathbf{x}^{\prime}\right\|_{2}^{2}}{2\ell^{2}}\right),\quad\mathbf{x},\mathbf{x}^{\prime}\in\mathcal{D}$ (A6)
GP-UCB algorithm.
$\mathbb{E}\left[f(\mathbf{x})\mid\mathbf{D}_{t}\right]=\mathbf{k}_{t}^{\top}(\mathbf{x})\left(\mathbf{K}_{t}+\sigma^{2}\mathbf{I}_{t}\right)^{-1}\mathbf{y}_{t},\operatorname{Var}\left[f(\mathbf{x})\mid\mathbf{D}_{t}\right]=k(\mathbf{x},\mathbf{x})-\mathbf{k}_{t}(\mathbf{x})^{\top}\left(\mathbf{K}_{t}+\sigma^{2}\mathbf{I}_{t}\right)^{-1}\mathbf{k}_{t}(\mathbf{x}),$ (A7)
The evaluation efficiency of GP-UCB.
whitehouse2023sublinear, wang2023regret showed that GP-UCB srinivas2012information regret upper bound for both expected cumulative regret and simple regret almost tightly matches the lower bound provided by scarlett2017lower. Specifically, for any constant $B>0$, the GP-UCB algorithm $\mathcal{A}_{\text{GPUCB}}$ achieves
$\sup_{\|f\|_{\mathcal{N}_{\Psi}(\mathcal{D})}\leq B}\mathbb{E}\left[\mathcal{R}_{C}(T;f,\mathcal{A}_{\text{GPUCB}})\right]=\begin{cases}O(T^{\frac{\nu+d}{2\nu+d}}),&\text{ for Matérn kernels, }\\ O(T^{\frac{1}{2}}\ln^{\frac{d}{2}}(T))&\text{ for SE kernels. }\end{cases}$ (A8)
$\sup_{\|f\|_{\mathcal{N}_{\Psi}(\mathcal{D})}\leq B}T(\epsilon,f,\mathcal{A}_{\mathrm{GPUCB}})=\begin{cases}O\left(\left(\frac{1}{\epsilon}\right)^{2+d/\nu}\right),&\text{for Matérn kernels,}\\ O\left(\frac{1}{\epsilon^{2}}\cdot\left(\log\frac{1}{\epsilon}\right)^{d+3}\right),&\text{for SE kernels.}\end{cases}$ (A9)
Appendix B Proofs
B.1 Auxiliary lemmas
Lemma 2 (Uniform first-order expansion).
Let $g=\nabla\mu(e)$ and $h=\nabla\sigma(e)$. There exist $C>0$ and $\varepsilon_{0}>0$ such that for all $\varepsilon\in(0,\varepsilon_{0}]$ and all $u\in\mathbb{R}^{d}$,
$\mu(e+\varepsilon u)$ $=\mu(e)+\varepsilon\,g^{\top}u\;+\;r_{\mu}(\varepsilon,u),$ $|r_{\mu}(\varepsilon,u)|$ $\leq C\varepsilon^{2},$ (A10)
$\sigma(e+\varepsilon u)$ $=\sigma(e)+\varepsilon\,h^{\top}u\;+\;r_{\sigma}(\varepsilon,u),$ $|r_{\sigma}(\varepsilon,u)|$ $\leq C\varepsilon^{2}.$ (A11)
Proof.
By the mean-value theorem with Lipschitz gradients, for some $\xi$ between $e$ and $e+\varepsilon u$, $\mu(e+\varepsilon u)=\mu(e)+\nabla\mu(\xi)^{\top}(\varepsilon u).$ Hence $\mu(e+\varepsilon u)=\mu(e)+\varepsilon g^{\top}u+\varepsilon(\nabla\mu(\xi)-\nabla\mu(e))^{\top}u,$ and $|\nabla\mu(\xi)-\nabla\mu(e)|\leq L_{\mu}\|\xi-e\|\leq L_{\mu}\varepsilon$, giving $|r_{\mu}|\leq L_{\mu}\varepsilon^{2}$. The proof for $\sigma$ is identical with $L_{\sigma}$. Taking $C:=\max\{L_{\mu},L_{\sigma}\}$ yields the stated bounds with the same $C$. ∎
Lemma 3 (Max-stability under bounded perturbations).
$g_{i_{f}}\geq f_{i_{f}}-\Delta\quad\text{and}\quad f_{i_{g}}\geq g_{i_{g}}-\Delta,$ (A12)
Proof.
Lemma 4 (Spherical cap coverage).
$\mathbb{P}\!\left(\max_{i\leq N}v^{\top}U_{i}\geq 1-\eta_{N}\right)\to 1.$ (A13)
Proof.
Lemma 5 (Near-maximum coupling in a thin band).
Let $M_{N}=\max_{i}\xi_{i}$ and define $v_{N}:=(g+M_{N}h)/\|g+M_{N}h\|$. Also, $S_{i}:=a_{i}+b_{i}M_{N}$, $Y_{i}^{(0)}:=a_{i}+b_{i}\xi_{i}$, $b_{i}:=\sigma(e)+\varepsilon h^{\top}U_{i}$. Fix any sequence $c_{N}\uparrow\infty$ with $c_{N}/q_{N}\to 0$ and set $\delta_{N}:=c_{N}/q_{N}$. Then with probability $\to 1$ there exists an index $j$ such that
$\xi_{j}\geq M_{N}-\delta_{N}\quad\text{and}\quad v_{N}^{\top}U_{j}\geq 1-\eta_{N},$ (A14)
whenever $\eta_{N}\downarrow 0$ satisfies
$N\,\mathbb{P}(\xi\geq M_{N}-\delta_{N})\,\rho\{u:v_{N}^{\top}u\geq 1-\eta_{N}\}\to\infty.$ (A15)
Moreover, for this $j$,
$S_{j}-Y_{j}^{(0)}=b_{j}(M_{N}-\xi_{j})\leq b_{\max}\,\delta_{N}.$ (A16)
Proof.
$E_{N}:=\{M_{N}\geq q_{N}-\delta_{N}\}.$ (A17)
$\mathbb{P}(\xi\geq t_{N})\;\geq\;\mathbb{P}(\xi\geq M_{N}-\delta_{N})\quad\text{on }E_{N}.$ (A18)
Let
$K_{N}:=\sum_{i=1}^{N}\mathbf{1}\{\xi_{i}\geq t_{N}\}\sim\mathrm{Binomial}\!\big(N,\,p_{N}\big),\qquad p_{N}:=\mathbb{P}(\xi\geq t_{N}).$ (A19)
A Chernoff bound gives
$\mathbb{P}\!\left(K_{N}<\tfrac{1}{2}Np_{N}\right)\;\leq\;\exp\!\left(-\tfrac{1}{8}Np_{N}\right).$ (A20)
Consider
$L_{N}\;:=\;\sum_{i=1}^{N}\mathbf{1}\{\xi_{i}\geq t_{N}\}\,\mathbf{1}\{v_{N}^{\top}U_{i}\geq 1-\eta_{N}\}.$ (A21)
[…]
$L_{N}\,\big|\,\{\xi_{i}\}\;\sim\;\mathrm{Binomial}\!\big(K_{N},\,p_{\mathrm{cap},N}\big),\qquad\mathbb{P}\big(L_{N}=0\,\big|\,\{\xi_{i}\}\big)\;=\;(1-p_{\mathrm{cap},N})^{K_{N}}\;\leq\;\exp\!\big(-K_{N}p_{\mathrm{cap},N}\big).$ (A22)
$\leq\mathbb{E}\!\left[\exp\!\big(-\tfrac{1}{2}Np_{N}\,p_{\mathrm{cap},N}\big)\right]\;+\;\exp\!\big(-\tfrac{1}{8}Np_{N}\big)\;+\;o(1).$ (A23)
On $E_{N}$, (A18) yields $p_{N}\geq\mathbb{P}(\xi\geq M_{N}-\delta_{N})$. By the lemma’s hypothesis,
Since $0\leq e^{-\frac{1}{2}Z_{N}}\leq 1$, for any fixed $T>0$,
Letting $T\to\infty$ gives $\mathbb{E}[e^{-\frac{1}{2}Z_{N}}]\to 0$. Moreover, $Z_{N}\leq Np_{N}$ since $p_{\mathrm{cap},N}\leq 1$. If $Np_{N}$ were bounded along a subsequence, then $Z_{N}$ would be bounded along that subsequence, contradicting $Z_{N}\xrightarrow{p}\infty$. Hence $Np_{N}\to\infty$, so $\exp(-\tfrac{1}{8}Np_{N})\to 0$. Therefore the second term in (A23) vanishes, and $\mathbb{P}(L_{N}=0)\to 0$.
B.2 Proof of Theorem 1
Proof.
By Lemma 2,
$Y_{i}$ $=\mu(e)+\varepsilon g^{\top}u_{i}+r_{\mu}(\varepsilon,u_{i})+\big(\sigma(e)+\varepsilon h^{\top}u_{i}+r_{\sigma}(\varepsilon,u_{i})\big)\xi_{i}$
$=a_{i}+b_{i}\xi_{i}+R_{i},$ (A24)
where $a_{i}:=\mu(e)+\varepsilon g^{\top}u_{i}$, $b_{i}:=\sigma(e)+\varepsilon h^{\top}u_{i}$, and $R_{i}:=r_{\mu}(\varepsilon,u_{i})+\xi_{i}r_{\sigma}(\varepsilon,u_{i}).$ Then $|R_{i}|\leq C\varepsilon^{2}(1+|\xi_{i}|)$ for all $i$. Define $S_{i}:=a_{i}+b_{i}M_{N}$, $b_{\max}:=\max_{1\leq i\leq N}|b_{i}|$, and $U_{i}:=u_{i}$.
Now let $Y_{i}^{(0)}:=a_{i}+b_{i}\xi_{i}$ and $i_{0}\in\arg\max_{i}Y_{i}^{(0)}$. From (A24),
By the bound on $R_{i}$ and $|\xi_{i}|\leq M_{N}:=\max_{j}\xi_{j}$,
$Y_{i^{\star}}^{(0)}\geq\max_{i}Y_{i}^{(0)}-2C\varepsilon^{2}(1+M_{N}).$ (A25)
Also,
where the last term is $\leq 0$ provided $b_{i}\geq 0$ for all $i$, which holds by Assumption A4 for $\varepsilon\leq\varepsilon_{0}:=\sigma(e)/(2\|h\|)$ when $\sigma(e)>0$ since $b_{i}=\sigma(e)+\varepsilon h^{\top}U_{i}\geq\sigma(e)-\varepsilon\|h\|\geq\tfrac{1}{2}\sigma(e)>0$. Therefore,
$\max_{i\leq N}S_{i}-S_{i^{\star}}\;\leq\;\varepsilon\|g+M_{N}h\|\,\eta_{N}\;+\;b_{\max}\delta_{N}\;+\;4C\varepsilon^{2}(1+M_{N}).$ (A26)
Set:
By Lemma 2,
$S_{i}\;=\;\mu(e)+M_{N}\sigma(e)+\varepsilon(g+M_{N}h)^{\top}u_{i}+\tilde{r}_{i},\qquad|\tilde{r}_{i}|\leq C\varepsilon^{2}(1+M_{N}).$
Hence for any $i$,
so
$\big|\,(S_{i}-S_{i^{\star}})-\varepsilon\|g+M_{N}h\|\,(v_{N}^{\top}U_{i}-v_{N}^{\top}\widehat{u}_{N})\,\big|\;\leq\;2C\varepsilon^{2}(1+M_{N}).$ (A27)
Taking $i$ that attains $\max_{i}S_{i}$ and using (A26) yields
$\max_{i\leq N}v_{N}^{\top}U_{i}-v_{N}^{\top}\widehat{u}_{N}\;\leq\;\eta_{N}\;+\;\frac{\Delta^{\prime}_{N}(\varepsilon)}{\varepsilon\|g+M_{N}h\|}.$ (A28)
Thus
[…]
Assume $\|h\|>0$ (the $\|h\|=0$ case is addressed in Remark 1). Using
$\|g+M_{N}h\|\;\geq\;M_{N}\|h\|-\|g\|$
and Lemma 1, $M_{N}\to\infty$ and $M_{N}\asymp q_{N}$. Also $b_{\max}\leq\sigma(e)+\varepsilon\|h\|+C\varepsilon^{2}$. Hence
$N\,e^{c_{N}}\,\rho\{u:v_{N}^{\top}u\geq 1-\eta_{N}\}\to\infty$
makes the RHS of (A29) equal to $o_{p}(1)+O(\varepsilon)$ for fixed $\varepsilon>0$; then let $\varepsilon\downarrow 0$.
With $N\to\infty$ first and fixed $\varepsilon>0$, (A29) implies $1-v_{N}^{\top}\widehat{u}_{N}\xrightarrow{p}0$. Then $\varepsilon\downarrow 0$ removes the $O(\varepsilon)$ residual, yielding $v_{N}^{\top}\widehat{u}_{N}\xrightarrow{p}1$. Since $M_{N}-q_{N}\to 0$ in probability by Lemma 1, the unit vectors $v_{N}=(g+M_{N}h)/\|g+M_{N}h\|$ and $\tilde{v}_{N}:=(g+q_{N}h)/\|g+q_{N}h\|$ satisfy $\|v_{N}-\tilde{v}_{N}\|\to 0$ in probability. Hence $\widehat{u}_{N}\xrightarrow{p}\tilde{v}_{N}$, which is the claimed gradient direction of $A_{\beta_{N}}$ with $\beta_{N}=q_{N}$.
By Lemma 2,
$A_{\beta_{N}}(e+\Delta e^{(N)})=A_{\beta_{N}}(e)+\varepsilon\nabla A_{\beta_{N}}(e)^{\top}\widehat{u}_{N}+O(\varepsilon^{2}).$
Write $\nabla A_{\beta_{N}}(e)=g+q_{N}h$. Using $\tilde{v}_{N}^{\top}\widehat{u}_{N}\xrightarrow{p}1$ gives $\nabla A_{\beta_{N}}(e)^{\top}\widehat{u}_{N}=\|\nabla A_{\beta_{N}}(e)\|\,\tilde{v}_{N}^{\top}\widehat{u}_{N}=\|\nabla A_{\beta_{N}}(e)\|(1-o_{p}(1)).$ Thus $A_{\beta_{N}}(e+\Delta e^{(N)})\geq A_{\beta_{N}}(e)+\varepsilon\|\nabla A_{\beta_{N}}(e)\|-o_{p}(\varepsilon),$ as claimed. ∎
Remark 1 (Edge cases).
If $\|h\|=0$, then $\nabla A_{\beta_{N}}(e)=g$ and the direction reduces to $\nabla\mu(e)$. The Bayesian Optimization setup does not allow $\sigma(e)=0$.
Appendix C Appendix for Section 7
C.1 Creative Briefs for the Scenarios
Below are the creative briefs for the eight ad campaigns. Each creative brief outlines the strategic and creative direction for each scenario. The selected campaigns cover a diverse range of products across distinct categories. Each campaign addresses a unique consumer need—whether related to sustainability, luxury, technology, or lifestyle—ensuring comprehensive coverage of various market segments. From environmentally-conscious millennials to affluent, experience-driven travelers, the approach accounts for the full spectrum of consumer interests. By focusing on different consumer needs, the structure enables a holistic marketing strategy that captures both niche and broad demands, providing a well-rounded framework for the study.
C.1.1 Scenario 1: GreenBite Plant-Based Burger
  • $\bullet$
    Product: "GreenBite," a new plant-based burger patty.
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
    Brand Personality & Tone: Positive, confident, and mouth-watering.
C.1.2 Scenario 2: AuraSonics X1 Earbuds
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
    Brand Personality & Tone: Minimalist, intelligent, and calm.
C.1.3 Scenario 3: Odyssey E-SUV
  • $\bullet$
    Product: The "Odyssey E-SUV," a new all-electric family SUV.
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
C.1.4 Scenario 4: Oasis Eco-Lodge
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
    Brand Personality & Tone: Elegant, peaceful, and understated.
C.1.5 Scenario 5: Momentum Digital Banking
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
    Brand Personality & Tone: Modern, empowering, and simple.
C.1.6 Scenario 6: MindGarden Meditation App
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
    Brand Personality & Tone: Gentle, approachable, and encouraging.
C.1.7 Scenario 7: Aeterno Watch
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
    Brand Personality & Tone: Elegant, timeless, and confident.
C.1.8 Scenario 8: SyncFlow B2B Software
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
  • $\bullet$
    Brand Personality & Tone: Professional, efficient, and innovative.
C.2 Prompts for Initial Ad Generation
Meta-prompt used for generating prompts for initial 64 ads generation CREATIVE BRIEF:
{creative_\_brief}
@@ C.2 Prompts for Initial Ad Generation
Based on this creative brief, generate a creative, structured, and descriptive prompt for a generative AI model (e.g., Imagen) that will produce a brand-aligned and scroll-stopping advertisement image suitable for an Instagram feed.
@@ 5. What to include:
Figure A1: Meta-prompt for generating the prompts for the initial 64 ads.
Worst-of-64 prompt for GreenBite (plant-based burger patty) ad scenario. **Key Message:** To visually communicate the core idea: "Experience the juicy, no-compromise satisfaction of a real burger, made better for you and the planet."
@@
Best-of-64 prompt for GreenBite (plant-based burger patty) ad scenario. **1. Key Message (The "Why"):** To visually capture the peak moment of guilt-free indulgence, proving that a plant-based burger can deliver an unbelievably juicy, satisfying, and no-compromise flavor experience identical to a classic beef burger. The text overlay will read: "The No-Compromise Burger."